# MECD+: Unlocking Event-Level Causal Graph Discovery for Video Reasoning

### v3

@@ IV-A VGCM: Video Granger Causality Model
where $\text{Enc}_{V}$ and $\text{Enc}_{T}$ represent the encoder module of video clips and captions, respectively. Dec is a multi-modal decoder that shares weights for both paths. Cat indicates the concatenate operation, and $[-]_{N-1}$ indicates the $(N-1)$-th slice at dimension 0. ${\bm{F}}^{p}$ and ${\bm{F}}^{m}_{k}$ are features after encoding, and ${\bm{O}}^{p}$ and ${\bm{O}}^{m}_{k}$ are the output features, which are then used for comparison of difference. Incorporating both visual and linguistic representations, the decoder conducts cross-modal reasoning and leverages the context from the unmasked premise events to posit a meaningful representation of the most likely explanatory result event.
@@ IV-B Causal Inference
Specifically, when $e_{k}$ is masked for the comparison in VGCM, the causal relations between $e_{k}$’s adjacent events and the last event are impacted, leading to a confounding of causal effects. Notably, for brevity, we only employ $e_{k}$’s previous one event $e_{k-1}$ and its subsequent one event $e_{k+1}$ for analysis, but the same analysis also applies to all the previous or subsequent events. To be specific, there exist two distinct kinds of confounding when $e_{k}$ is absent: 1) Causal effects of $e_{k-1}$ to $e_{N}$ may be lost, as its connection to $e_{N}$ is built upon $e_{k}$, (green path in Fig. 6 (a1)). 2) Causal effects of $e_{k+1}$ to $e_{N}$ may be redundant, as $e_{k}$ must be a necessary prior of its causality, (red path in Fig. 6 (a1)).
@@ IV-C Context Chain Reasoning
Specifically, during the training phase, we adopt a strategy of concurrently masking multiple events, $e_{k}=\{e_{k0},...,e_{ki}\}$, (where $0<i<(n-1)$), to enhance the model’s consideration of contextual events during reasoning. The ground truth causal relation is represented by the labels of the multiple events $e_{k}=\{e_{k0},...,e_{ki}\}$ that are masked simultaneously, which are combined using an OR operation: $r_{k}=\bigvee_{j=0}^{i}r_{kj}$.
@@ IV-D Non-regressive Complete Graph Reasoning
Through the Event Causality Test introduced in the previous sections, we can infer the causal chains between all premise events and the final event $e_{N}$. To further infer the complete event-level causal graph during the inference phase, a straightforward method treats each event as a result event and conducts Event Causality Test using a regressive approach [67], ultimately synthesizing a complete causal graph. However, this regressive approach is highly inefficient because inferring a causal graph for a video with $N$ events requires the model to perform forward propagation $(N-1)(N-2)/2$ times. This quadratic time complexity is suboptimal, especially given that our dataset contains an average of more than five events per video. To tackle this, we employ the following efficient non-regressive complete graph reasoning method during inference.
@@
where $k$ equals $i$, as $k$ iterates from $1$ to $N-1$ during inference of causal relations with $e_{N}$, no additional forward propagation is required. Our approximation method specifically involves utilizing the features ${\bm{O}}^{p}$ and ${\bm{O}}^{m}_{i}$ without additional masking of $e_{j}$, following the approach outlined in  [67]. The ground truth result event $e_{j}$ is encoded in the same manner as $e_{i}$, yielding the encoded feature ${\bm{F}}_{j}=\text{Enc}_{V}(\Phi_{pre}(v_{j}))$ and the output feature ${\bm{O}}_{j}=\text{Dec}(\texttt{{Cat}}({\bm{F}}_{j},\text{Enc}_{C}({\bm{C}}_{j}))$. Following a similar mask-based reasoning as Eq. 2, after reasoning, the output causal relation $\hat{r}_{[i:j]}$ can be represented by:
@@
As a result, just like inferring causal chains only related to the result event $e_{N}$, where inference requires just $(N-1)$ forward propagation, this non-regressive complete graph reasoning method does not require additional time-consuming backward reasoning, therefore equips the VGCM with rapid inference speed.
@@ V-A Implementation details
Figure 8: An Example of the Evaluation of SHD. The figure illustrates that the prediction of the complete causal graph on the right contains two missing edges and two extra edges. Consequently, the total number of incorrectly predicted causal relation edges is four, leading to a SHD value of 4.
@@
All the experiments of VGCM are conducted on 1 NVIDIA A6000 GPU. We train VGCM for about 8 hours. VideoLLMs fine-tuning is conducted on 8 A6000 GPUs.
Comparisons: We mainly compared our model with baseline model Videobert [68] and widely used multi-modal base models such as CLIP-L [72], SIGLIP [73], and the most representative video reasoning model VAR [17]. Besides, we also conduct experiments on powerful LLM, including Mixtral-8x22B-Instruct [74], GPT-4 [50], Gemini-Pro [49], and so on. Additionally, ImageLLMs and VideoLLMs utilized for comparison include widely accepted GPT4-o [50], Gemini-Pro [49], VideoLLaVA [75], VideoChat2 [76] and so on.
In the setting of few-shot learning (In-Context Learning), Large Language Models (LLMs) and ImageLLMs are evaluated, following the methodologies employed in natural language processing for causal discovery tasks [77, 78, 79]. To more comprehensively assess the capacity for causal reasoning and to leverage the benefits of open-source models, the performance of VideoLLMs and all multi-modal base models is assessed under a strong fine-tuning paradigm. For all VideoLLMs, the vision-language projector was fully fine-tuned using Lora on the training set of the MECD dataset. Similarly, for multi-modal base models, the vision and text encoders are frozen, and a causal relation classifier, composed of the MLP layers and sigmoid function, is fully trained on the training set of the MECD dataset.
TABLE II: Main Results. VGCM reaches SOTA performance on causal chains and complete causal graph reasoning. “wo CI” indicates without causal inference, while “wo CR” indicates without context chain reasoning.
@@
In-context Llama-3-70B [80] 5.10 67.81 52.33 60.00
GLM-4 (0520)  [81] 4.86 67.50 52.47 60.08
DeepSeek-Coder [82] 5.11 67.27 55.01 61.97
@@
Metrics: Our model has two output formats: one that outputs causal chains related to the result event, focusing on the model’s multi-cause reasoning ability, and another that outputs a complete causal graph of all events, ultimately reflecting the generalization ability in causal reasoning. Therefore, we designed two metrics to evaluate these aspects respectively.
@@ V-B Main Results
TABLE III: Inference Speed. Our VGCM is 3-6 times faster than all VideoLLMs while slightly slower than the baseline.
@@
MiniGPT-4 [84] 4.72
a) Our Performances. As shown in Tab. II, our VGCM without causal inference reaches an accuracy of 66.95%, demonstrating basic video causal reasoning capabilities. Furtherly, the complete VGCM outperforming the SOTA Proprietary LLM GPT-4 [50], ImageLLM GPT4-o [50], open-source basic multi-modal model SIGLIP [73], VideoLLM VideoChat2 [76] by 6.40%, 5.77%, 6.43%, and 2.70%, respectively.
@@
Additionally, our VGCM further highlights its advantages in terms of complete causal graph reasoning, particularly without additional data supervision annotations. VGCM achieves performance with only an average of 3.94 false causal relations for a complete causal graph. An example of the output complete causal graph is visualized in Fig. 9.
As illustrated in Tab. LABEL:speed, we have assessed the inference speed of various models, with our VGCM achieving a swift 0.76 seconds per sample. The proposed method incurs an overhead of only 8.57% over the Videobert [68] baseline. It is noteworthy that our inference speed is 3 to 6 times faster than that of all VideoLLMs. Additionally, our VGCM is also faster than the CLIP-B and CLIP-L [72] model, further demonstrating the efficiency of our model’s causal reasoning.
@@
c) VideoLLM Performances. As shown in Tab. II, under the paradigm of In-Context Learning, models such as VideoChat2 [76] and PLLaVA [86] have demonstrated superior performance, likely due to the inclusion of Next-QA [8] and CLEVRER [9] datasets in their pre-training data, which share characteristics with the causal aspects of the MECD dataset. However, we found that VideoLLMs do not significantly surpass LLMs under the same framework of In-Context Learning, which may be attributed to the parameter limitations of VideoLLMs and the limitations of multi-modal In-Context Learning [90, 91]. Furthermore, under the paradigm of Fine-tuning, the performances are enhanced, though they still fall short of the performance of our VGCM.
Moreover, the performance of VideoLLMs suggests a reduction in the disparity between the Pos and Neg metrics, which may be attributed to the mitigation of hallucinations and the decrease in the influence of caption ambiguity due to adequate visual information. Furthermore, the enhancement in the Pos metric for VideoChat2 [76] could be attributed to the utilization of spatiotemporal cues to infer causal relations [76]. As shown in Fig. 12, after utilizing spatiotemporal cues, compared to the same encoders with MLP structure model EVA-CLIP [92], VideoChat2 [76] successfully discovered the long-term causal relations between the initial reason event “paying in cash” and the result event “having a quarrel”, reasoning leads to the ultimate source of results.
Figure 12: Comparison between Output Causal Chains of EVA-CLIP and VideoChat2. VideoChat2 could be attributed to the utilization of spatiotemporal cues to infer causal relations.
@@
d) Human Performances. As shown in Tab. II, the average performance of ten volunteers reaches 87.19%. When only one modality of information serves as the input, we observe that annotators exhibit higher accuracy when visual information is the sole input. We posit that this may be attributed to limitations in caption quality, such as the difficulty in judging certain scene transitions or pronouns that can obscure the identification of the person. An example of caption confusion is shown in Fig. 13, however, this problem can be released with additional visual input.
@@
$\text{VideoLLM}^{*}$ Minigpt4-Video [85] 61.61 -4.11
Minigpt-4 [84] 61.69 -1.02
@@ V-C Ablation Study
The highest performance was achieved when Videobert [68] was employed as the encoder. Pre-training methods that rely on self-supervision have been demonstrated to significantly enhance the performance of downstream visual tasks, including video frame prediction [96]. Similarly, the superior performance of the Videobert encoder may be attributed to the fact that the self-supervised pre-training of Videobert [68] augments the masked contrastive learning inherent in the VGCM model, then further improves the causal reasoning ability of the VGCM model.
@@
d) Input Modalities. As the MECD task employs both video input and corresponding captions, we aim to evaluate the degree of reliance on these two modalities. Typically, each event in our dataset consists of a textual input with an average of 13.5 words caption and a visual input of 50 frames.
@@
0.10.30.40.866666666686868687070707072727272Non-maskedText-maskedVision-masked
To investigate the influence of the text modality, we employ a masking strategy for the input caption of the premise event, gradually increasing the masking ratio from 10% to 80%. The results presented in Fig. 15 indicate that our VGCM does not rely on the textual modality input; VGCM can also conduct causal discovery for videos without any captions.
In contrast, the experimental results suggest a more obvious performance decrease towards less visual modality input in the causality discovery task, as shown in Fig. 15. However, even with 80% masking of either modality, the results consistently outperform fine-tuned VideoLLaVA [75], underscoring the robust causal discovery capability of VGCM.
@@
1) Reasoning Abilities with LLMs Versus MLPs. We compare the performances between VideoLLMs and corresponding basic multi-modal models in Tab. IX. We found the understanding capabilities of LLMs significantly enhance the causal discovery of multi-modal foundation models versus simple MLP layers. Although visual and textual information are encoded by the same encoder, the models are equipped with enhanced reasoning capabilities through a more sophisticated 7B language model. Especially in LanguageBind [97] encoders, VideoLLaVA [75] with LanguageBind encoders outperforms the original models by up to 7.29% on accuracy.
@@
2) Input Modalities. We compare the performances between VideoLLMs with all modalities input and only vision input as shown in Tab. X. When additional text information is input into VideoLLMs, it serves to supplement the limitations of visual information and emphasizes key details. However, ambiguous text cues can exacerbate illusions, which in turn leads to a decline in the Pos metric that contrasts with the Acc trend. The modest increase in the overall indicator of causal reasoning ability (+1.43) underscores the necessity for the model to prioritize visual information and exhibit a high Visual-Textual Information Ratio, thereby maintaining a greater proportion of visual to textual data [98, 46].
@@
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.911116666666668686868707070707272727274747474PercentageTOP-1 Accuracy (%)Random Flipped DataData Volume
@@ VI Conclusion
We proposed a novel task, multi-event video causal discovery (MECD), which focuses on event-level complete causal graph reasoning. To support this task, we developed the MECD dataset, consisting of 1,438 long-term daily-life videos featuring intricate causal relationships. Additionally, we proposed the first model for video causal discovery, the Video Granger Causality Model (VGCM), which is based on the principles of the Event Granger Test and incorporates advanced causal inference techniques to address issues such as illusions and confounding factors. VGCM employs robust and efficient reasoning mechanisms, including context chain reasoning and non-regressive complete graph reasoning. Our VGCM outperformed GPT-4o and VideoChat2 by 5.77% and 2.70%, respectively, highlighting its superior reasoning capabilities.

### v4

@@ IV-A VGCM: Video Granger Causality Model
where $\text{Enc}_{V}$ and $\text{Enc}_{C}$ represent the encoder module of video clips and captions, respectively. Dec is a multi-modal decoder that shares weights for both paths. Cat indicates the concatenate operation, and $_{N-1}$ indicates the (N1)(N-1)-th slice at dimension 0. ${\bm{F}}^{p}$ and ${\bm{F}}^{m}_{k}$ are features after encoding, and ${\bm{O}}^{p}$ and ${\bm{O}}^{m}_{k}$ are the output features, which are then used for comparison of difference. Incorporating both visual and linguistic representations, the decoder conducts cross-modal reasoning and leverages the context from the unmasked premise events to posit a meaningful representation of the most likely explanatory result event.
@@ IV-B Causal Inference
Specifically, when $e_{k}$ is masked for the comparison in VGCM, the causal relations between $e_{k}$’s adjacent events and the last event are impacted, leading to a confounding of causal effects. Notably, for brevity, we only employ $e_{k}$’s previous one event $e_{k-1}$ and its subsequent one event $e_{k+1}$ for analysis, but the same analysis also applies to all the previous or subsequent events. To be specific, there exist two distinct kinds of confounding when $e_{k}$ is absent: 1) Causal effects of $e_{k-1}$ to $e_{N}$ may be lost, as its connection to $e_{N}$ is built upon $e_{k}$, (green path in Fig. 6 (a1)). 2) Causal effects of $e_{k+1}$ to $e_{N}$ may be redundant, as $e_{k}$ may be a necessary prior of $e_{k+1}$’s connection with $e_{N}$, (red path in Fig. 6 (a1)).
@@ IV-C Context Chain Reasoning
Specifically, during the training phase, we adopt a strategy of concurrently masking multiple events, $e_{k}=\{e_{k0},...,e_{ki}\}$, (where 0<i<(n1)0<i<(n-1)), to enhance the model’s consideration of contextual events during reasoning. The ground truth causal relation is represented by the labels of the multiple events $e_{k}=\{e_{k0},...,e_{ki}\}$ that are masked simultaneously, which are combined using an OR operation: $r_{k}=\bigvee_{j=0}^{i}r_{kj}$.
@@ IV-D Non-regressive Complete Graph Reasoning
Through the Event Causality Test introduced in the previous sections, we can infer the causal chains between all premise events and the final event $e_{N}$. To further infer the complete event-level causal graph during the inference phase, a straightforward method treats each event as a result event and conducts Event Causality Test using a regressive approach [70], ultimately synthesizing a complete causal graph. However, this regressive approach is highly inefficient because inferring a causal graph for a video with $N$ events requires the model to perform forward propagation (N1)(N2)/2(N-1)(N-2)/2 times. This quadratic time complexity is suboptimal, especially given that our dataset contains an average of more than five events per video. To tackle this, we employ the following efficient non-regressive complete graph reasoning method during inference.
@@
where $k$ equals $i$, as $k$ iterates from $1$ to N1N-1 during inference of causal relations with $e_{N}$, no additional forward propagation is required. Our approximation method specifically involves utilizing the features ${\bm{O}}^{p}$ and ${\bm{O}}^{m}_{i}$ without additional masking of $e_{j}$, following the approach outlined in  [70]. The ground truth result event $e_{j}$ is encoded in the same manner as $e_{i}$, yielding the encoded feature ${\bm{F}}_{j}=\text{Enc}_{V}(\Phi_{pre}(v_{j}))$ and the output feature ${\bm{O}}_{j}=\text{Dec}(\texttt{{Cat}}({\bm{F}}_{j},\text{Enc}_{C}({\bm{C}}_{j}))$. Following a similar mask-based reasoning as Eq. 2, after reasoning, the output causal relation $\hat{r}_{[i:j]}$ can be represented by:
@@
As a result, just like inferring causal chains only related to the result event $e_{N}$, where inference requires just (N1)(N-1) forward propagation, this non-regressive complete graph reasoning method equips the VGCM with rapid inference speed.
@@ V-A Implementation details
Figure 8: An Example of the Evaluation of SHD. The figure illustrates that the prediction of the complete causal graph on the right contains two missing edges and one extra edge. Consequently, the total number of incorrectly predicted causal relation edges (SHD value) is 3.
@@
Comparisons: We compare VGCM with baseline [71] and multi-modal base models such as CLIP-L [75], and a representative video reasoning model VAR [15]. Besides, we also conduct experiments on powerful LLM, including Mixtral-8x22B-Instruct [76], GPT-4 [53], Gemini-Pro [52], and so on. Additionally, ImageLLMs and VideoLLMs utilized for comparison include widely accepted GPT4-o [53], Gemini-Pro [52], VideoLLaVA [77], VideoChat2 [78], Long-VideoLLM LongVA [79], Qwen2.5-VL [80], and so on.
In the setting of few-shot learning (In-Context Learning), LLMs and ImageLLMs are evaluated, following the principle utilized in causal discovery tasks [81, 82, 83] in NLP. To more comprehensively assess the capacity for causal reasoning and to leverage the benefits of open-source models, the performance of VideoLLMs and all multi-modal base models is assessed under a strong fine-tuning paradigm. For all VideoLLMs, the vision-language projector was fully fine-tuned using Lora on the training set of the MECD dataset. Similarly, for multi-modal base models, the vision and text encoders are frozen, and a causal relation classifier, composed of the MLP layers and sigmoid function, is fully tuned.
TABLE II: Main Results. VGCM reaches SOTA performance on causal chains and complete causal graph discovery. “w/o CR” indicates without context chain reasoning.
@@
In-context DeepSeek-Coder [84] 5.11 67.27 55.01 61.97
@@
Metrics: Our model has two output formats: one that outputs causal chains related to the result event, focusing on the model’s multi-cause reasoning ability (towards causal chains discovery task), and another that outputs a complete causal graph of all events (towards complete causal graph discovery task), ultimately reflecting the generalization ability in causal reasoning. Therefore, we designed two metrics to evaluate these aspects respectively.
@@ V-B Main Results
TABLE III: Inference Speed. VGCM is 3-6 times faster than all VideoLLMs while slightly slower than the baseline. w/o NR indicates without non-regressive complete graph reasoning.
@@
a) Our Performances. As shown in Tab. II, our VGCM outperforming the SOTA Proprietary LLM GPT-4 [53], ImageLLM GPT4-o [53], open-source basic multi-modal model SIGLIP [89], VideoLLM VideoChat2 [78] by 6.40%, 5.77%, 6.43%, and 2.70%, respectively.
@@
Additionally, our VGCM further highlights its advantages in terms of complete causal graph reasoning, particularly without additional data supervision annotations. VGCM achieves performance with only an average of 3.94 false causal relations for a complete causal graph. An example is shown in Fig. 9.
As illustrated in Tab. LABEL:speed, we have assessed the inference speed of various models, with our VGCM achieving a swift 0.76 seconds per sample. VGCM with non-regressive complete graph reasoning incurs an overhead of only 8.57% over the Videobert [71] baseline. VGCM’s inference speed is 3 to 6 times faster than that of all VideoLLMs. This shows that VGCM with the non-regressive approach achieves both precise and efficient reasoning capabilities.
@@
c) VideoLLM Performances. As shown in Tab. II, under the paradigm of In-Context Learning, models such as VideoChat2 [78] and PLLaVA [88] have demonstrated superior performance, likely due to the inclusion of Next-QA [6] and CLEVRER [7] datasets in their pre-training data. Furthermore, under the paradigm of Fine-tuning, the performances are enhanced, though they still fall short of the performance of our VGCM. Moreover, under more frames of input, Long-VideoLLMs (LongVA, Qwen2.5-VL) do not outperform general VideoLLMs. However, their relatively better SHD performance indicates that long-sequence perception enhances the ability to reason about overall causal relations.
Besides, the performance of VideoLLMs suggests a reduction in the disparity between the Pos and Neg metrics, which may be attributed to the mitigation of hallucinations and the decrease in the influence of caption ambiguity due to adequate visual information. Furthermore, the enhancement in the Pos metric for VideoChat2 [78] could be attributed to the utilization of spatiotemporal cues to infer causal relations [78]. As shown in Fig. 12, after utilizing spatiotemporal cues, compared to the same encoders with MLP structure model EVA-CLIP [94], VideoChat2 [78] successfully discovered the long-term causal relations between the initial reason event “paying in cash” and the result event “having a quarrel”, reasoning leads to the ultimate source of results.
The reasons for the performance gap between VideoLLMs and VGCM are summarized here: (1) VGCM is tailored for causal reasoning, while general VideoLLMs lack causal understanding due to insufficient exposure during pretraining; (2) VGCM models bidirectional dependencies using the Granger Causality framework, enabling robust causal inference, while the autoregressive nature of VideoLLMs enforces an autoregressive bias that limits their ability to capture global event relations.
Figure 12: Comparison between Output Causal Chains of EVA-CLIP and VideoChat2. VideoChat2 could be attributed to utilizing spatiotemporal cues to infer causal relations.
@@
d) Human Performances. As shown in Tab. II, the average performance of ten volunteers reaches 87.19%. When only inputting one modality of information, higher accuracy is achieved when visual information is the sole input. We posit that this may be attributed to ambiguous pronouns can affect the judgment of their referents. An example of caption confusion is shown in Fig. 13, however, this problem can be released with additional visual input.
@@
$\text{VideoLLM}^{*}$ Minigpt-4 [86] 61.69 -1.02
@@ V-C Ablation Study
The highest performance was achieved when Videobert [71] was employed as the encoder, the core reason is as follows. The mask-prediction training approach in Videobert [71] is particularly effective for learning contextual relations, making it well-suited for subsequent causal reasoning guided by Granger Causality. The principle of Granger Causality is to assess whether preceding event information improves the prediction of future outcomes—a concept that aligns with Videobert’s objective of modeling contextual dependencies.
@@
d) Input Modalities. Typically, the input of VGCM consists of a textual input with an average of 13.5 words and a visual input of 50 frames. To investigate the influence of the text modality, we employ a masking strategy for the input caption of the premise event, gradually increasing the masking ratio from 10% to 80%. The results presented in Fig. 15 indicate that VGCM does not rely on the textual modality input.
@@
0.10.30.40.86666686870707272Non-maskedText-maskedVision-masked
In contrast, the experimental results suggest a more obvious performance decrease towards less visual modality input in the causality discovery task, as shown in Fig. 15.
@@
1) Reasoning Abilities with LLMs Versus MLPs. We compare the performances between VideoLLMs and corresponding basic multi-modal models in Tab. IX. We found the understanding capabilities of LLMs significantly enhance the causal discovery of multi-modal foundation models versus simple MLP layers. Although visual and textual information are encoded by the same encoder, the models are equipped with enhanced reasoning capabilities through a more sophisticated 7B language model.
@@
2) Input Modalities of VideoLLMs. We compare the performances between VideoLLMs with all modalities input and only vision input as shown in Tab. X. When additional text information is input into VideoLLMs, it serves to supplement the limitations of visual information and emphasizes key details. However, ambiguous text cues can exacerbate illusions, which in turn leads to a decline in the Pos metric that contrasts with the Acc trend. The modest increase in the overall indicator of causal reasoning ability (+1.43) underscores a high Visual-Textual Information Ratio of the MECD dataset [78].
@@
00.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.91166666868707072727474PercentageTOP-1 Accuracy (%)Random Flipped DataData Volume
@@ VI Conclusion
We proposed a novel task, multi-event video causal discovery (MECD), which focuses on event-level complete causal graph reasoning. To support this task, we developed the MECD dataset, consisting of 1,438 long-term daily-life videos featuring intricate causal relations. Additionally, we proposed the first model for video causal discovery, the Video Granger Causality Model (VGCM), which is based on the principles of the Event Granger Test and incorporates advanced causal inference techniques to address issues such as illusions and confounding factors. VGCM employs robust and efficient reasoning mechanisms, including context chain reasoning and non-regressive complete graph reasoning. Our VGCM outperformed GPT-4o and VideoChat2 by 5.77% and 2.70%, respectively, highlighting its superior reasoning capabilities.